series of 13, 18, 16, 21, 19
Answer by checkley77(12844)

series of 13, 18, 16, 21, 19
The sequence is +5, -2, +5, -2, +5, -2
13+5=18-2=16+5=21-2=19+5=[24]-2=[22]+5=[27]